OpenStudy (anonymous):

-3x+3y=3, 3x+y=9 solve by using the elimination method

OpenStudy (anonymous):

-3x+3y=3, 3x+y=9 solve by using the elimination method

OpenStudy (unklerhaukus):

\[-3x+3y=3\]\[ \quad3x+y=9\] add the two equations

OpenStudy (anonymous):

I got the answer 3=9=12 what do i do next?

OpenStudy (unklerhaukus):

well you must have made a mistake three does not equal nine which does not equal twelve

OpenStudy (anonymous):

3+9=12 sorry typing error.

OpenStudy (unklerhaukus):

\[−3x+3y=3\]\[\quad3x+y=9\] adding the equations \[(−3x+3y)+(3x+y)=(3)+(9)\]now simplify and solve for \(y\)

OpenStudy (anonymous):

looks like @UnkleRhaukus had this one :) @UnkleRhaukus is AWESOME at math!!!! wootwoot for @UnkleRhaukus !!! :D

OpenStudy (anonymous):

iheartfood you're an awesome tutor as well so don't be putting yourseld down lol ;)

OpenStudy (anonymous):

haha geee thanks @juniormints :) lol :) I'm just sayinngggg @UnkleRhaukus is waaayyy better hahaa @UnkleRhaukus helps ME all the time :) hahaa but glad u think that :)

OpenStudy (unklerhaukus):

have you found \(y\) yet?